OpenStudy (arbyj):

Someone help me solve; 4(4y+1)=4+4y ? (please show work)

OpenStudy (anonymous):

So first you distribute the left side of the queation: 16y + 4 = 4 + 4y Next, you will move the variables on one side: 16y-4y=4-4 Now Solve: 12y=0 y=0

OpenStudy (arbyj):

Ok thank you so much @Safira97

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.